# Small Molecule Drug Discovery Through Deep Learning: Progress, Challenges, and Opportunities

### v1

@@ 3.4 Molecule Property Prediction
| Diffusion | Protein |
[1]Li et al. (2022a); [2]Huang et al. (2022); [3]Zhu et al. (2023); [4]Jo et al. (2022); [5]Lee et al. (2023b); [6]Guan et al. (2023);
Notations: and are the generator and the discriminator, denotes the variant SMILES and is the value function. and are the graph and geometry per fragment. Meanwhile, and are linear drift coefficients, and are scalar diffusion coefficients, and , are reverse-time standard Wiener processes. represents the OOD condition and is the parameter controlling the OOD generative process. Finally, denotes the set of atoms of the binding site, signifies the set of fragments per molecule, and is the prior covariance matrix. indicates that the -th molecule atom corresponds to the -th prior. are noise parameters.
@@
Various representations, such as SMILES, graph-based representations, and one-dimensional (1D) descriptions (e.g., molecular fingerprints and atom types), are used to encode molecular structures. The representation directly impacts prediction model performance. Tables 3 and 4 provide a detailed overview of the various methods with or without pre-training and their technical routes in predicting different molecular properties. Since different methods vary in dataset partitioning, evaluation metrics, random seeds, and implementation details, we selected representative methods and provided detailed descriptions of molecular representation and feature encoding techniques without comparing performance. The recent methods emphasize pre-training techniques (e.g., self-supervised and multi-modal contrastive learning, and 3D-based denoising tasks), which use large datasets to model and optimize the molecular feature space, fine-tuning it according to specific datasets. This pre-training technique can learn generic molecular feature representations from large amounts of labeled (or unlabeled) 3D, textual, or image data, thereby demonstrating excellent generalization performance on small-scale labeled data. In addition, MPP also encounters OOD challenges, especially when dataset splits are based on molecular scaffolds or similarity features, leading to poor predictions for structurally distinct samples.
@@
[1]Wu et al. (2024c); [2]Jin et al. (2023); [3]Huang et al. (2024); [4]Fu et al. (2022); [5]Zhu et al. (2024).
Notations: represents the generative conditions, represents -th atomic tags for the molecules’ SMILES strings, is the Transformer framework and Emb is the standard input for the transformer, encompassing both word and positional embeddings. Meanwhile, denotes the node weight vector and is the graph neural network whose parameters are , Finally, is the constantly updated model training dataset, and represents the batch size.
@@
[1]The therapeutics data commons dataset; [2]The collection of benchmark datasets for learning with graphs; [3]The open graph benchmark; [4]The benchmark for molecular machine learning; [5]Two molecular dynamics benchmarks for evaluating force fields; [6]Two molecular dynamics datasets; [7]Two DFT-based molecular datasets for computational chemistry; [8]A 3D geometries benchmark; [9]A benchmark dataset for equivariant many-body interaction operations;[10]A dataset of 20 million conformations; [11]A dataset with high-confidence data; [12]A Caenorhabditis elegans database; [13]The human DTI dataset; [14]A database of useful (docking) decoys-Enhanced; [15]The G protein-coupled receptor focused dataset;[16]A DTA and side effect dataset; [17]The biomedical network dataset; [18]A drug combination synergy database; [19]The gene&genome database;[20]A drug-side effect association database;[21]A DDI network dataset; [22]The biological knowledge with gene-drug-disease triples;[23]A multi-source dataset; [24]The cancer cell line encyclopedia; [25]The NCI-60 human tumor cell lines screen.
Notations: Quantum mechanics (Qm.); Biophysics (Bio.); Physical chemistry (Phys. Chem.); Physiology (Physio.).
3.5 Molecular Generation
The chemical space is vast, with over possible compound structures Zhu et al. (2023), making traditional methods slow and labor-intensive. Computational MG methods have thus become essential to accelerate molecule discovery with desired properties.
Initially, MG methods were largely unconditional random generation Huang et al. (2022); Jo et al. (2022), which failed to meet practical requirements. As a result, conditional MG methods have emerged as the mainstream approach. Key trends include 1) Property-conditioned generation Zhu et al. (2023), which targets molecules with specific properties. 2) Molecular (sub)structure-conditioned generation, which uses predefined structural elements for design. 3) Target-conditioned generation Guan et al. (2023), which focuses on molecules with high binding affinity to biological targets. 4) Phenotype-conditioned generation, which guides molecular design towards desired biological outcomes. Diffusion models have recently become dominant in MG Lee et al. (2023b), as they iteratively refine molecular structures from noise, offering more flexibility and integration with conditional frameworks for targeted molecular design.
@@ 3.6 Molecular Optimization
MO approaches can be broadly classified into two categories: generative models and search-based methods. Generative models like VAE, generative adversarial networks (GAN), flow-based methods, and diffusion models optimize molecules by manipulating their latent representations or refining structures iteratively from noise Huang et al. (2024); Xiong et al. (2024). In contrast, search-based methods, which use evolutionary algorithms Fu et al. (2022), navigate the chemical space more directly. Additionally, auto-regressive models construct molecules sequentially, allowing precise control over the optimization process Wu et al. (2024c). While early methods focused on latent space optimization, search-based approaches have gained popularity due to their superior interpretability and stability. Evolutionary algorithms evolve and eliminate molecules from a candidate set, while scoring functions modify molecules until the optimization objective is achieved. Other search-based methods map molecules to a high-dimensional solution space for optimization.
@@
4 Datasets
Small molecule drug discovery datasets form the foundation for research in this field. Typically, these datasets include information on the small molecules’ characteristics such as their chemical structures, biological activities, pharmacokinetic properties, and toxicity, spanning multiple stages from target identification and lead compound screening to drug optimization. To facilitate research and application, these datasets are categorized based on their data structures and task types, with corresponding volumes provided. Since these datasets are suitable for molecular generation and optimization tasks, Table 8 does not specifically list the datasets related to these two tasks. Additionally, due to the large number and variety of datasets, we provide the access URLs for each dataset without citing the references individually.
We present several commonly used benchmark and non-benchmark datasets. As shown in Table 8, the TDC datasets cover various tasks during small molecule drug discovery and are relatively mainstream. Additionally, QM9 and ZINC from TUDataset are molecular datasets that include a wide range of properties, such as quantum chemical properties, and are widely used in property prediction and conditional MG and MO tasks. Furthermore, the Open Graph Benchmark (OGB) and MoleculeNet databases include multiple molecular property datasets, which are commonly used for validating the performance of various molecular graph representation methods. For instance, MoleculeNet is a widely used benchmark dataset for molecular studies, encompassing four major categories: quantum mechanics, physical chemistry, biophysics, and physiology. Since the DRP, DDI, and DTI/DTA tasks investigate complex relationships between drugs and other entities (such as targets, cell lines, and drugs), the datasets involved are diverse and complex, and are thereby categorized under ”Others.” In addition, most small molecule datasets are regularly updated and expanded, and the dataset statistics are accurate as of February 4, 2025.
5 Challenges and Opportunities
@@
In the small molecule drug discovery field, achieving model interpretability through correlation (i.e., identifying which molecular features are strongly associated with the predicted outcomes) is relatively straightforward. This can be accomplished using traditional feature importance evaluation methods, such as SHapley Additive exPlanations or model-based attention mechanisms, which reveal the features that are strongly correlated with drug responses or drug–target interactions Fu et al. (2022); Zhu et al. (2024). For example, methods such as SubCDR Liu and Zhang (2023) decomposed molecules into multiple subcomponents and quantified their contributions in the prediction, providing explanations and helping us understand the model’s decision-making process. However, causal interpretability is more complex and challenging. Explaining the complex interactions between multiple entities requires understanding a molecule’s influence on the activities and responses of another through specific mechanisms, necessitating more complex models and methods. Moreover, causal inference typically relies on multi-gene regulatory networks that involve interactions among various factors, requiring a deep understanding of biological background knowledge. Therefore, achieving causal interpretability may require more specific techniques and experimental methods. For instance, constructing gene regulatory networks, integrating interdisciplinary datasets, or combining these with wet-lab experimental validation could offer feasible solutions to address this challenge.
@@
To improve model generalizability in OOD scenarios, several strategies may be useful. First, data augmentation and generative models, such as VAEs and GANs, can be used to simulate various ”drug+X” combinations, introducing the model to a broader range of potential scenarios during training and enhancing its adaptability to new environments. Second, transfer learning can leverage knowledge from tasks with known drugs or targets and apply it to new ones, assisting the model to address unseen drugs or X. Finally, domain adaptation techniques have recently demonstrated the ability to reduce distributional discrepancies between training and testing datasets Li et al. (2024b); He et al. (2022) by learning the mapping relationships between the source and target domains, further enhancing the model’s predictive ability on new data.

### v2

@@ III-D Molecule Property Prediction
| Diffusion | Protein | ||
| Diffusion | Property |
Notations: denotes a set of pairwise distances between atoms. and are the graph and geometry per fragment. and are coupling layers which process atoms and bonds respectively. denotes the distribution of atomic positions. models the distribution of inter-atomic distances given the graph and models the distribution of conformations given the distances . and are the generator and the discriminator, denotes the variant SMILES and is the value function. denotes the fake data generated by generator in GAN framework. Meanwhile, and are linear drift coefficients, and are scalar diffusion coefficients, and , are reverse-time standard Wiener processes. For a 3D molecule, it can be represented as point clouds , where is the atom coordinates matrix and is the node feature matrix. represents the OOD condition and is the parameter controlling the OOD generative process. Finally, denotes the set of atoms of the binding site, signifies the set of fragments per molecule, and is the prior covariance matrix. indicates that the -th molecule atom corresponds to the -th prior. are noise parameters. represents the generative conditions and denotes the scale of conditional guidance. is a parameterized mean function consisting of a deblurring network and is the small amount of standard deviation for noise. denotes the 3D structure of fragment coordinates.
@@
Various representations, such as SMILES, graph-based representations, and one-dimensional (1D) descriptions (e.g., molecular fingerprints and atom types), are used to encode molecular structures. The representation directly impacts prediction model performance. Tables III and IV provide a detailed overview of the various methods with or without pre-training and their technical routes in predicting different molecular properties. Since different methods vary in dataset partitioning, evaluation metrics, random seeds, and implementation details, we selected representative methods and provided detailed descriptions of molecular representation and feature encoding techniques without comparing performance. The recent methods emphasize pre-training techniques (e.g., self-supervised and multi-modal contrastive learning, and 3D-based denoising tasks), which use large datasets to model and optimize the molecular feature space, fine-tuning it according to specific datasets. This pre-training technique can learn generic molecular feature representations from large amounts of labeled (or unlabeled) 3D, textual, or image data, thereby demonstrating excellent generalization performance on small-scale labeled data. In addition, MPP also encounters OOD challenges, especially when dataset splits are based on molecular scaffolds or similarity features, leading to poor predictions for structurally distinct samples.
@@
| Diffusion | Protein | ||
| Diffusion | Cell Line | ||
| Search | Property | ||
| Search | Property | ||
| Search | Property | ||
| Search | Property | ||
@@
Notations: denote two-part latent representations of junction tree and graph structure. represents the generative conditions, represents -th atomic tags for the molecules’ SMILES strings, is the Transformer framework and Emb is the standard input for the transformer, encompassing both word and positional embeddings. is the reference arm. denotes the element-wise product, ”kn” and ”ukn” are the abbreviations for ”known” and ”unknown”, respectively. The noise , is a conditional control strength parameter (), and indicates unconditional generation. is the the molecular graph editing model and is the constantly updated model training dataset. are probability distributions with different operations on molecules. denotes the number of child nodes in the search procedure, is the average reward of node , is the number of times node is accessed, and is the iterative epochs. is the reward vector dimension, signifies an exploration parameter used to balance exploration and utilization, denotes the set of nodes based on the Pareto optimality selection. Meanwhile, denotes the node weight vector and is the graph neural network whose parameters are , Finally, is the constantly updated model training dataset, and represents the batch size. is the element in the sequence of a molecule and is the maximum length of the sequence.
@@
[1]Therapeutics data commons dataset; [2]The collection of benchmark datasets for learning with graphs; [3]The open graph benchmark; [4]The benchmark for molecular machine learning; [5]Molecular dynamics benchmark for evaluating force fields; [6]Two molecular dynamics datasets; [7]Two DFT-based molecular datasets for computational chemistry; [8]A 3D geometries benchmark; [9]A benchmark dataset for equivariant many-body interaction operations;[10]A dataset of 20 million conformations; [11]A dataset with high-confidence data; [12]A Caenorhabditis elegans database; [13]The human DTI dataset; [14]A database of useful (docking) decoys-Enhanced; [15]The G protein-coupled receptor focused dataset;[16]A DTA and side effect dataset; [17]The biomedical network dataset; [18]A drug combination synergy database; [19]The gene&genome database;[20]A drug-side effect association database;[21]A DDI network dataset; [22]The biological knowledge with gene-drug-disease triples;[23]A multi-source dataset; [24]The cancer cell line encyclopedia; [25]The NCI-60 human tumor cell lines screen.
Notations: Quantum mechanics (Qm.); Biophysics (Bio.); Physical chemistry (Phys. Chem.); Physiology (Physio.). The dataset statistics are accurate as of February 4, 2025.
III-E Molecular Generation
The chemical space is vast, with over possible compound structures [65], making traditional methods slow and labor-intensive. Computational MG methods have thus become essential to accelerate molecule discovery with desired properties.
Table VI provides a comprehensive summary. Initially, MG methods were largely unconditional random generation [63, 69], which failed to meet practical requirements. As a result, conditional MG methods have emerged as the mainstream approach. Key trends include 1) Property-conditioned generation [65], which targets molecules with specific properties. 2) Molecular (sub)structure-conditioned generation, which uses predefined structural elements for design. 3) Target-conditioned generation [72], which focuses on molecules with high binding affinity to biological targets. 4) Phenotype-conditioned generation, which guides molecular design towards desired biological outcomes. Diffusion models have recently become dominant in MG [71], as they iteratively refine molecular structures from noise, offering more flexibility and integration with conditional frameworks for targeted molecular design.
@@ III-F Molecular Optimization
Table VII provides a comprehensive summary. MO approaches can be broadly classified into two categories: generative models and search-based methods. Generative models like VAE, generative adversarial networks (GAN), flow-based methods, and diffusion models optimize molecules by manipulating their latent representations or refining structures iteratively from noise [79, 87]. In contrast, search-based methods, which use evolutionary algorithms [84], navigate the chemical space more directly. Additionally, auto-regressive models construct molecules sequentially, allowing precise control over the optimization process [76]. While early methods focused on latent space optimization, search-based approaches have gained popularity due to their superior interpretability and stability. Evolutionary algorithms evolve and eliminate molecules from a candidate set, while scoring functions modify molecules until the optimization objective is achieved. Other search-based methods map molecules to a high-dimensional solution space for optimization.
@@
IV Datasets
Small molecule drug discovery datasets form the foundation for research in this field. Typically, these datasets include information on the small molecules’ characteristics such as their chemical structures, biological activities, pharmacokinetic properties, and toxicity, spanning multiple stages from target identification and lead compound screening to drug optimization. To facilitate research and application, these datasets are categorized based on their data structures and task types, with corresponding volumes provided. Since these datasets are suitable for molecular generation and optimization tasks, Table VIII does not specifically list the datasets related to these two tasks. Additionally, due to the large number and variety of datasets, we provide the access URLs for each dataset without citing the references individually.
We present several commonly used benchmark and non-benchmark datasets. As shown in Table VIII, the TDC datasets cover various tasks during small molecule drug discovery and are relatively mainstream. Additionally, QM9 and ZINC from TUDataset are molecular datasets that include a wide range of properties, such as quantum chemical properties, and are widely used in property prediction and conditional MG and MO tasks. Furthermore, the Open Graph Benchmark (OGB) and MoleculeNet databases include multiple molecular property datasets, which are commonly used for validating the performance of various molecular graph representation methods. For instance, MoleculeNet is a widely used benchmark dataset for molecular studies, encompassing four major categories: quantum mechanics, physical chemistry, biophysics, and physiology. Since the DRP, DDI, and DTI/DTA tasks investigate complex relationships between drugs and other entities (such as targets, cell lines, and drugs), the datasets involved are diverse and complex, and are thereby categorized under ”Others.”
V Challenges and Opportunities
@@
In the small molecule drug discovery field, achieving model interpretability through correlation (i.e., identifying which molecular features are strongly associated with the predicted outcomes) is relatively straightforward. This can be accomplished using traditional feature importance evaluation methods, such as SHapley Additive exPlanations or model-based attention mechanisms, which reveal the features that are strongly correlated with drug responses or drug–target interactions [84, 85]. For example, methods such as SubCDR [16] decomposed molecules into multiple subcomponents and quantified their contributions in the prediction, providing explanations and helping us understand the model’s decision-making process. However, causal interpretability is more complex and challenging. Explaining the complex interactions between multiple entities requires understanding a molecule’s influence on the activities and responses of another through specific mechanisms, necessitating more complex models and methods. Moreover, causal inference typically relies on multi-gene regulatory networks that involve interactions among various factors, requiring a deep understanding of biological background knowledge. Therefore, achieving causal interpretability may require more specific techniques and experimental methods. For instance, constructing gene regulatory networks, integrating interdisciplinary datasets, or combining these with wet-lab experimental validation could offer feasible solutions to address this challenge.
@@
To improve model generalizability in OOD scenarios, several strategies may be useful. First, data augmentation and generative models, such as VAEs and GANs, can be used to simulate various ”drug+X” combinations, introducing the model to a broader range of potential scenarios during training and enhancing its adaptability to new environments. Second, transfer learning can leverage knowledge from tasks with known drugs or targets and apply it to new ones, assisting the model to address unseen drugs or X. Finally, domain adaptation techniques have recently demonstrated the ability to reduce distributional discrepancies between training and testing datasets [21, 22] by learning the mapping relationships between the source and target domains, further enhancing the model’s predictive ability on new data.